OpenStudy (anonymous):

PLEASE HELP LAST ASSIGNMENT IN CLASS MUST FINISH IN NEXT 2 HRS!!!!

OpenStudy (anonymous):

I know it's a lot but it's not even an 1/8 of the entire assignment

OpenStudy (anonymous):

i got the answer for the - Let c be any constant number. Which of the following will always be perpendicular to −3x + y = −1? so don't worry about it

OpenStudy (anonymous):

I got the venn diagram on too the- A news agent conducted a survey among business magazine subscribers of a town and found that there was circulation of only three business magazines. He made the Venn diagram below to show the number of subscribers to each of the three magazines: Board Review, Investor Journal, and Business Talk.

OpenStudy (anonymous):

@zepp can you PLEASE help me ;(

OpenStudy (anonymous):

i got the- Which ratio represents the area of the smaller rectangle compared to the area of the larger rectangle? (Figure not drawn to scale) one too

OpenStudy (ash2326):

[7.03] A film set designer is using white and colored tiles in a pattern to create paths of different lengths. If x is the length of the path in feet, the number of colored tiles needed to make it is calculated using the rule (6x2y3)(4xy3) where y represents the number of white tiles. Which simplified expression represents the number of colored tiles used for a path of length x feet? 24x2y6 24x3y6 24x2y9 24x3y9 @Ashley99 did you try this one?

OpenStudy (anonymous):

i got the volume one- In a kitchen there are four containers that can hold different quantities of water, as shown in the figure below.

OpenStudy (anonymous):

yeh it's 24x3y6 right?

OpenStudy (anonymous):

sorry forgot to put that i did that one :)

OpenStudy (anonymous):

the next one is 5p3q right?

OpenStudy (ash2326):

Yeah you're correc

OpenStudy (ash2326):

correct for the first

OpenStudy (ash2326):

Second is also correct, good work:D

OpenStudy (anonymous):

the second is 5p^3q and the 3rd is 14 right????

OpenStudy (anonymous):

k what about 5?(i'm working on 4 now)

OpenStudy (ash2326):

second and third are right, good work @Ashley99

OpenStudy (anonymous):

thnx 4 is -2 can you help w/ 5 plz

OpenStudy (anonymous):

5 is last one on that doc.

OpenStudy (anonymous):

(i did all the other ones)

OpenStudy (ash2326):

sure, if a line passes through two points say (x1, y1) and (x2, y2) then its equation is given as \[\frac{y-y2}{x-x2}=\frac{y2-y1}{x2-x1}\] just plugin the values and you'll get the equation (x1, y1)=(1, 1) (x2, y2)=(-2, 13)

OpenStudy (anonymous):

please continue w/ #5 explanation

OpenStudy (ash2326):

Okay, I'll plugin the values \[\frac{y-13}{x-(-2)}=\frac{13-1}{-2-1}\] \[\frac{y-13}{x+2}=\frac{12}{-3}\] \[\frac{y-13}{x+2}=-4\] \[{y-13}={(x+2)}\times-4\] \[y-13=-4x+8\] Can you simplify it now?

OpenStudy (anonymous):

@ash2326 yes i can but can you continue and show me hoe so know my anser is right? plz ;(

OpenStudy (ash2326):

It's just two steps now, try it, if you get stuck I'll help you

OpenStudy (anonymous):

-8 on each side so it becomes y-24=-4x

OpenStudy (anonymous):

y-6=x right?

OpenStudy (ash2326):

We have \[y-13=-4x+8\] add 13 to both the sides

OpenStudy (ash2326):

sorry it;s \[y-13=-4x-8\]

OpenStudy (anonymous):

sorry wrong variable

OpenStudy (anonymous):

y=-4x+5

OpenStudy (ash2326):

now add 4x to both the sides

OpenStudy (anonymous):

4x+y=5

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

And #6=1756 right??

OpenStudy (anonymous):

well really it's 1756.3 but same difference

OpenStudy (ash2326):

We have \[x-0.3x=1351\] you'll find x from here but we need to find the total sales in both the months so it's actually x+1351

OpenStudy (anonymous):

this one is not a multi-choice one plz help

OpenStudy (ash2326):

How many visitors did the website record in both months combined? so the answer is x+1351

OpenStudy (ash2326):

plugin x=1756

OpenStudy (anonymous):

no it was 504 for the 1st month 1351 for the second and 1756 total

OpenStudy (anonymous):

1855 sorry :)

OpenStudy (ash2326):

no. of visitors in second month=1351 this is 30 % less than the first month's Let the first month's sale be x so \[30 \%\ of\ x=\frac{30}{100}x=0.3x\] so \[x-0.3x=1351\] find x, which comes out as 1756 so total no. of visitors in two months=x+1351=1756+1351

OpenStudy (ash2326):

sorry* let the no. of visitors in first month be x

OpenStudy (anonymous):

total of 3107 i'm looking at my work it seems i had it right originally but then i started going in circles :)

OpenStudy (ash2326):

no problem , did you understand it?

OpenStudy (anonymous):

yeh i'll message you if if any more questions i need help w/ i still have 47 more to solve ;(

OpenStudy (ash2326):

oops, try as many as you could, if you don't get any or if you get stuck. Post it here:)

OpenStudy (anonymous):

thanks i already have a headache so i might need more help than normal but i will try to limit the posted questions

OpenStudy (anonymous):

@ash2326 thanks for everything :)

OpenStudy (ash2326):

@Ashley99 you're welcome, take rest and then study:) You'll get more correct then :D

OpenStudy (anonymous):

if i had the time i would thanks anyway :)